### v2

# Hypergeometric orthogonal polynomials of Jacobi type

Abstract.
Motivated by the theory of hypergeometric orthogonal polynomials, we consider quasi-orthogonal polynomial families – those that are orthogonal with respect to a non-degenerate bilinear form defined by a linear functional – in which the ratio of successive coefficients is given by a rational function $f(u,s)$ which is polynomial in $u$. We call this a family of Jacobi type.
Our main result is that, up to rescaling and renormalization, there are only five families of Jacobi type. These are the classical families of Jacobi, Laguerre and Bessel polynomials, and two more one parameter families $E_{n}^{(c)},F_{n}^{(c)}$. Each family arises as a specialization of some hypergeometric series. The last two families can also be expressed through Lommel polynomials, and they are orthogonal with respect to a positive measure on $\mathbb{R}$ for $c>0$ and $c>-1$ respectively.
We also consider the more general rational families, i.e. quasi-orthogonal families in which the ratio $f(u,s)$ of successive coefficients is allowed to be rational in $u$ as well.
1991 Mathematics Subject Classification:
33C45, 33D45
1. Introduction
1.1. Orthogonal Polynomials
Let ${\mathcal{P}}=(P_{0},P_{1},\ldots,P_{n},\ldots)$ be a family of polynomials in $\mathbb{C}[z]$, where $P_{i}=P_{i}(z)$ is a polynomial of degree $i$.
We will say that ${\mathcal{P}}$ is a quasi-orthogonal family if there exists a linear functional $M:\mathbb{C}[z]\to\mathbb{C}$ such that the matrix with entries $Q_{ij}=M(P_{i}P_{j})$ is an invertible diagonal matrix – in other words, we have $Q_{ij}=0$ if $i\neq j$ and $Q_{ij}\neq 0$ if $i=j$.
The functional $M$ is usually called a moment functional. It is easily seen that $M$ completely determines ${\mathcal{P}}$ up to a renormalization of the form $P_{i}\mapsto c_{i}P_{i},\;c_{i}\in\mathbb{C}^{\times}$. In particular there is a unique family for which each $P_{i}$ is monic. Conversely, the functional $M$ is uniquely determined by the quasi-orthogonal family $\mathcal{P}$ and the normalization $M(1)=1$.
We say that a quasi-orthogonal family ${\mathcal{P}}$ is orthogonal if all the polynomials $P_{i}$ are defined over $\mathbb{R}$ and there exists a moment functional $M$ such that all the diagonal entries $Q_{ii}$ are real and strictly positive. It is known that in this case the moment functional $M$ is defined by integration with some positive measure $\mu$ on $\mathbb{R}$ that has an infinite support, and ${\mathcal{P}}$ is, up to normalization, the standard family of orthogonal polynomials defined by this measure (see e.g. [Chi78, Ch.II] or [Ism09, Ch.2]).
There are many classical (quasi-)orthogonal families of polynomials. These families play an important role in many areas of Mathematics, Physics and Engineering (see §1.3 for some references).
The characteristic property of these classical families is that they can be described in terms of hypergeometric functions. There is a general idea to axiomatize the notion of a hypergeometric (quasi-)orthogonal family and try to classify all such families.
One such attempt is presented by Askey-Wilson scheme (see §1.3 below). We will see that this scheme misses some natural families. The difficulty is that the notion of a hypergeometric family is not easy to define formally, and so it is difficult to axiomatize.
In this paper we propose an axiomatic description and give a classification of some class of hypergeometric (quasi-)orthogonal families of polynomials that we call families of Jacobi type.
Let us write the polynomials in our family ${\mathcal{P}}$ explicitly in terms of their coefficients $P_{n}=\sum_{k=0}^{n}c(n,k)z^{k}$. We extend the domain of definition of the coefficients $c(n,k)$ to $\mathbb{Z}^{2}$ by setting $c(n,k)=0$ outside of the range $0\leq k\leq n$. Roughly speaking, we will say that a quasi-orthogonal family $\mathcal{P}$ is of rational type if there exists a rational function $f(u,s)$ such that $c(n,k+1)\equiv f(n,k)c(n,k)$, We will say that a rational type family is of Jacobi type if $f$ depends polynomially on the first variable $u$. Since $f$ can have poles we need to be more careful.
Definition 1.1.
Let ${\mathcal{P}}$ be a quasi-orthogonal family of polynomials with coefficients $c(n,k)$. We say that ${\mathcal{P}}$ is of rational type if there exist relatively prime polynomials $N(u,s)$ and $D(u,s)$ such that
(1) $c(n,k+1)D(n,k)=c(n,k)N(n,k)\quad\text{ for all }n,k\in\mathbb{Z}.$
We say that ${\mathcal{P}}$ is of Jacobi type if it is of rational type, and $D$ does not depend on $s$, i.e. $D\in\mathbb{C}[u]$.
A priori the polynomials $N$ and $D$ could have a finite number of common zeroes. This would mean that we do not have conditions on a finite number of coefficients $c(n,k)$. Nevertheless, in Lemma 6.1 below we show that the rational function $N/D$ determines ${\mathcal{P}}$ uniquely up to renormalization.
1.2. Main result
To formulate our main result, we recall the definition of the hypergeometric series
(2) $_{p}F_{q}(\underline{a};\underline{b};z)=\sum_{k\geq 0}\frac{(a_{1})_{k}\cdots(a_{p})_{k}}{(b_{1})_{k}\cdots(b_{q})_{k}k!}z^{k},\quad(c)_{k}:=c(c+1)\cdots(c+k-1),$
and note that if $a_{1}=-n$ then the infinite series (2) truncates to give a polynomial of degree $\leq n$.
Theorem A.
Any family ${\mathcal{P}}$ of Jacobi type arises by a rescaling $P_{n}(z)\mapsto P_{n}(cz)$ and/or a renormalization $P_{n}(z)\mapsto c_{n}P_{n}(z)$ from one of the five hypergeometric series listed below. These families are quasi-orthogonal (“quasi”) and orthogonal (“orth”) for the indicated parameters.
  1. (a)
    Jacobi: ${}_{2}F_{1}(-n,n+a;b;z)$;  quasi: $a,b-1,a-b\notin\mathbb{Z}_{<0}$; orth: $a>b-1,\,b>0,\,a,b\in\mathbb{R}$;
  2. (b)
    Bessel: ${}_{2}F_{0}(-n,n+a;;z)$;  quasi: $a\notin\mathbb{Z}_{<0}$; orth: $a>-1,\,a\in\mathbb{R}$;
  3. (c)
    Laguerre: ${}_{1}F_{1}(-n;b;z)$;  quasi: $b\notin\mathbb{Z}_{\leq 0}$; orth: $b>0,\,b\in\mathbb{R}$;
  4. (d)
    $E_{n}^{(c)}$: ${}_{4}F_{1}(-n,-n-c+1,n+c,n+1;1/2;z)$; quasi: $c\notin\mathbb{Z}_{\leq 0}$; orth: $c>0,\,c\in\mathbb{R}$;
  5. (e)
    $F_{n}^{(c)}$: ${}_{4}F_{1}(-n,-n-c+1,n+c+1,n+2;3/2;z)$; quasi: $c\notin\mathbb{Z}_{\leq 0}$; orth: $c>-1,c\neq 0,\,c\in\mathbb{R}$.
The Jacobi, Bessel, and Laguerre families are, of course, classical. As for the other two families, it turns out that if one combines $E_{n}^{(c)}$ and $F_{n}^{(c)}$ into a single family $G_{n}^{(c)}$ defined as follows
(3) $G_{2n}^{(c)}=(-1)^{n}E_{n}^{(c)}(-z^{2}),\quad G_{2n+1}^{(c)}=(-1)^{n}zF_{n}^{(c)}(-z^{2}),$
then the polynomials $G_{n}^{(c)}$ satisfy the remarkably simple recursion $G_{n+1}^{(c)}=2(c+n)zG_{n}^{(c)}-G_{n-1}^{(c)}.$ It follows readily that the $G_{n}^{(c)}$ coincide with the Lommel polynomials $h_{n}^{(c)}$, which arise in the study of zeros of Bessel functions ([Wat44, §9.6], [Ism09, §6.5]). The polynomials $h_{n}^{(c)}$ themselves are alternately even and odd, and thus have many vanishing coefficients. In particular they are not a Jacobi type family. In this respect they are similar to the Hermite polynomials, which are not of Jacobi type, but which can be obtained as in (3) from two Laguerre families.
Our proof of Theorem A is purely algebraic. The main point of our argument is that starting with a rational family ${\mathcal{P}}$ we construct some algebraic object - a module $M$ that encodes this family. This module has much more structure than the family ${\mathcal{P}}$, so it is much easier to analyze.
More precisely, consider the field $K=\mathbb{C}(u,s)$ of bivariate rational functions. Denote by $A$ the subalgebra of $\operatorname{End}_{\mathbb{C}}(K)$ generated by multiplication operators and the shift operators $S^{\pm}:s\mapsto s\pm 1,\;U^{\pm}:u\mapsto u\pm 1$. Starting with a family ${\mathcal{P}}$, we construct an $A$-module $M(\mathcal{P})$. The condition that the family ${\mathcal{P}}$ is quasi-orthogonal can be written as a 3-term recursion relation. This gives very strong restrictions on the module $M(\mathcal{P})$. For example, as our first step we prove that this module is one dimensional over the field $K$. After this we use a general classification of one dimensional modules to fix the shape of this module. This general approach allowed us to prove a theorem that considerably restricts the possibilities for families of rational type.
Our second main result is that for any family $\mathcal{P}$ of rational type, the module $M(\mathcal{P})$ is isomorphic to $M(\mathcal{Q})$, where $\mathcal{Q}$ is either a family of Jacobi type, or is given by $Q_{n}=\,_{4}F_{1}(-n,-n+d,n+a,n+c;b;z)$ for some scalars $a,b,c,d\in\mathbb{C}$. We actually prove a slightly stronger theorem, that can be formulated in the following way.
Theorem B.
For any family ${\mathcal{P}}$ of rational type there exists polynomial $g\in\mathbb{C}[u,s]$ such that ${\mathcal{P}}$ arises by a rescaling and/or a renormalization from the family $P_{n}=g(n,z\partial_{z})Q_{n}$, where either
  1. (a)
    $Q_{n}$ is a Jacobi, Bessel, or Laguerre family
  2. (b)
    There exist scalars $a,b,c,d\in\mathbb{C}$ such that
    (4) $Q_{n}=\,_{4}F_{1}(-n,-n+d,n+a,n+c;b;z)$
Remark 1.2.
We expect that in case (b), the set of quadriples $(a,b,c,d)$ that define a rational family is quite restricted. It is quite possible that we can always find $g$ such that $Q_{n}$ is given by $a=1$, $b=1/2$ and $d=1-c$, i.e. that $Q_{n}=E_{n}^{(c)}$. In particular, if $P_{n}=F_{n}^{(c)}$ we have $g=(s+1/2)(u+s+c)(u+s+1)$.
In Appendix B we give several examples of more complicated $g$, for which $Q_{n}$ is not of Jacobi type. In them $Q_{n}$ is either a Jacobi family, or $E_{n}^{(c)}$. Let us describe here one such family $R_{n}^{\lambda}$. Let $Q_{n}$ be the Jacobi family with $a=1,b=3/2$, and $g(u,s):=\lambda(u+1/2)+s+1/2$, where $\lambda\notin\{-1,1\}$. Then
(5) $R^{\lambda}_{n}=g(u,z\partial_{z})Q_{n}=c_{n}\,{}_{3}F_{2}(-n,n+1,\lambda n+\frac{\lambda+3}{2};3/2,\lambda n+\frac{\lambda+1}{2};z),$
where $c_{n}=(\lambda+1)(n+1/2)$. This family is of rational type, is orthogonal and satisfies a third order differential equation. For further details on this family we refer the reader to §B.1.1.
1.3. Background, related results, future directions
The theory of orthogonal polynomials on the real line is a classical subject with applications to many areas of mathematics. We refer the reader to [Wat44, WW21] for the classical theory and to [Ism09, KWKS10, KLS10, AAR99, GR04, Gro78] for more recent developments.
It is instructive to compare our result with that of Bochner [Boc29], who classified second order differential operators with polynomial eigenfunctions. The classic Sturm-Liouville theorem provides sufficient conditions for such eigenfunctions to constitute an orthogonal family, and in Bochner’s case this yields the Jacobi, Laguerre, Bessel, and Hermite polynomials. Note however that the eigenfunction condition is a very strong requirement, which has the immediate effect of restricting the set of possible families to lie in a three dimensional parameter space, and which makes the classification problem quite a bit easier [Ism09]. In contrast, the classification result in Theorem A is obtained in the infinite dimensional parameter space corresponding to the coefficients of $f(u,s)$ in $\mathbb{C}(s)[u]$. Thus it is remarkable that there are only two additional orthogonal families in this general setting, both of which come from hypergeometric series of the form (2).
There are other ways of obtaining orthogonal families from hypergeometric functions and from their $q$-analogs, the basic hypergeometric functions ${}_{r}\phi_{s}$. The well-known Askey scheme [KLS10] is a hierarchical organization of some 44 families which arise in this manner. Most of these families can be obtained by limiting procedures applied to the Askey-Wilson polynomial, which is a ${}_{4}\phi_{3}$ family that depends on four parameters $a,b,c,d$ in addition to $q$. Ismail [Ism09] describes a generalization of Bochner’s theorem to second order $q$-difference operators and shows orthogonal families arising in this way are limits of Askey-Wilson polynomials. More recently Verde-Star [V-S21] has shown that almost all of the families in the Askey scheme belong to a class of families characterized by three recursive sequences. Still more recently, Koornwinder [Koo22] has obtained abstract characterization theorems for the Askey scheme within the Verde-Star class.
Just as in the Bochner setting, the assumptions of Ismail and Verde-Star immediately restrict the possible families to lie in a finite dimensional parameter space – 6 dimensions for Ismail and 11 dimensions for Verde-Star. It seems to us however that one can hope for a classification of hypergeometric and $q$-hypergeometric orthogonal families in a much greater generality, and we plan to return to this question in subsequent work.
We note also that the Askey-Wilson polynomials and their various limits can be understood in terms of a certain representation of a double affine Hecke algebra (DAHA) of rank 1 [Sah99, Sah00]. Thus one may ask whether the new polynomials can be understood in terms of a generalization of the DAHA theory.
1.4. Structure of the paper and ideas of the proofs
In §2 we give some necessary preliminaries, including a classification result of one-dimensional $A$-modules from [Ore30] (see Theorem 2.3 below), and Favard’s theorem, which states that a family of monic polynomials is quasi-orthogonal if and only if there exist sequences $\alpha_{n}$ and $\beta_{n}$ of complex numbers such that
(6) $zP_{n}=P_{n+1}+\alpha_{n}P_{n}+\beta_{n}{P_{n-1}}\text{ for all }n\geq 1;$
In §3 we construct an $A$-module from the coefficient function $c(n,k)$ and prove that it is one-dimensional. We first state that $f$ determines the family $P_{n}$, and that the sequences $\alpha$ and $\beta$ coincide with rational functions, except possibly for finitely many $n$. Then we consider the space of all $\mathbb{C}$-valued functions on $\mathbb{Z}^{2}$ as a module over $\mathbb{C}[u,s]$, and extend scalars to obtain a vector space over the field of rational functions $K=\mathbb{C}(u,s)$. This vector space has a natural structure of an $A$-module. Then we consider the submodule $M$ generated by the image of $c(n,k)$, that we denote by $w$. By construction we have $S^{-1}w=(S^{-1}f)w$, and (6) gives a linear relation on $S^{-1}w,\,Uw,\,w,$ and $U^{-1}w$ over the subfield $\mathbb{C}(u)$. Using this relation and the commutation $US=SU$ we prove in Theorem 3.5 below that $\dim_{K}M=1$.
In §4 we use the explicit realization from Theorem 2.3 to almost classify all pairs $M,v$, where $M$ is a one-dimensional $A$-module, and $v$ is a vector satisfying a relation of the form (6). Namely, we realize $v$ as a function
$\Phi=\exp(cu+ds)g(s,u)\prod\Gamma(k_{i}u+l_{i}s+c_{i}),$
where $g\in K=\mathbb{C}(u,s)$, satisfying
(7) $S^{-1}\Phi=U\Phi+\alpha\Phi+\beta U^{-1}\Phi.$
For every $i$, $\Phi$ has an infinite family of poles in the lines $k_{i}u+l_{i}s+c_{i}=n$ for $n\in\mathbb{Z}_{\geq 0}$. Any of the poles of one of the terms in (7) has to be a pole of at least one other term. From this we deduce that $k_{i},l_{i}\in\{-1,0,1\}$. Further analyzing the poles we obtain that there exist polynomials $p,q,w\in\mathbb{C}[t]$ and $g\in\mathbb{C}[u,s]$ such that
(8) $f=\frac{S\Phi}{\Phi}=\frac{q(u+s)p(s-u)}{w(s)}\cdot\frac{Sg}{g}$
In the rest of the section we use elementary algebraic considerations to deduce from (7) strong restrictions on $p,q$, and $w$ (see Corollaries 4.7 and 4.9). In particular, we show that $\dim p,\dim q,\dim w\leq 2$.
In §5 we prove Theorems A and Theorem B. We first deduce Theorem B from (8) and Corollary 4.7. Then we show that the condition that in families of Jacobi type $f$ depends polynomially on $u$ implies that in (8) we can assume that $g=1$. Finally, we deduce Theorem A from Theorem B and Corollary 4.9.
In §6 we prove further properties of families of Jacobi type. In particular, we explicitly compute the coefficients $\alpha$ and $\beta$ in (6) for each of the cases.
In §7 we investigate the two new families $E_{n}^{(c)}$ and $F_{n}^{(c)}$, express them through Lommel polynomials, and find the discrete measures on the real line with respect to which they are orthogonal. We also give the 4th order differential equations that they satisfy.
In Appendix A we prove the technical lemmas from §3. In particular, we prove that $f$ determines the family $P_{n}$, and that the sequences $\alpha_{n}$ and $\beta_{n}$ coincide with rational functions, except possibly for finitely many $n$.
In Appendix B we construct a two-parameter family ${\mathcal{P}}^{(c,\lambda)}$ of rational type that interpolates between the families $E^{(c+1)}$ and $F^{(c)}$, and is not of Jacobi type for $\lambda\notin\{0,1\}$. We also construct a two-parameter family of rational type that interpolates between two families of Jacobi polynomials.
2. Preliminaries
2.1. Favard’s theorem
Let $\{P_{n}\}_{n=0}^{\infty}$ be a family of monic polynomials.
Theorem 2.1 (see e.g.[Chi78, §4]).
  1. (i)
    The family $\{P_{n}\}$ is quasi-orthogonal if and only if there exist (unique) sequences $\alpha_{n}$ and $\beta_{n}$ of complex numbers, and $c_{0}\in\mathbb{C}$ such that
    (9) $zP_{n}=P_{n+1}+\alpha_{n}P_{n}+\beta_{n}{P_{n-1}}\,\,\forall n\in\mathbb{Z}_{\geq 1}$
    and $\beta_{n}\neq 0$ for all $n\in\mathbb{Z}_{\geq 1}$.
  2. (ii)
    A quasi-orthogonal family $\{P_{n}\}$ is orthogonal if and only if all the polynomials $P_{n}$ have real coefficients, and for all $n\in\mathbb{Z}_{\geq 1}$ we have $\alpha_{n},\beta_{n}\in\mathbb{R}$, and $\beta_{n}>0$.
2.2. One-dimensional $A$-modules
Recall the notation $K=\mathbb{C}(u,s)$ and $A:=$ the subalgebra of $\operatorname{End}_{\mathbb{C}}(K)$ generated by $K$ and the shift operators $U^{\pm 1}:u\mapsto u\pm 1;\,S^{\pm 1}:s\mapsto s\pm 1$. The $S^{\pm 1},U^{\pm 1}$ act on $K$ by $(U^{\pm 1}g)(u,s):=g(u\pm 1,s)$ and $(S^{\pm 1}g)(u,s):=g(u,s\pm 1)$. We say that an $A$-module is one-dimensional if it is one-dimensional as a vector space over $K$.
For every one-dimensional module $M$ and every non-zero vector $v\in M$ there exist unique $f,h\in K$ such that $Sv=fv$ and $Uv=hv$. It is easy to see that
(10) $Uf/f=Sh/h$
and that if we replace $v$ by $gv$ for some non-zero $g\in K$, the pair $(f,h)$ will change to $(fSg/g,hUg/g)$. We will call such pairs equivalent. Conversely, it is easy to see that every pair $(f,h)\in K^{\times}\times K^{\times}$ satisfying (10) defines a one-dimensional $A$-module. The set of isomorphism classes of one-dimensional $A$-modules forms a group under tensor product over $K$, where the $U$ and the $S$ act on a product by acting on both factors. This corresponds to elementwise products of pairs $(f,h)$. This group was computed in [Ore30]. To formulate this result, note that the field of meromorphic functions in $u,s$ has a natural $A$-module structure.
Definition 2.2.
A meromorphic function $\eta$ in two variables $u,s$ is of $\gamma$ type if it is a product of the form
(11) $\exp(au+bs)\prod_{i=1}^{n}\Gamma(k_{i}u+l_{i}s+c_{i}),\quad a,b,c_{i}\in\mathbb{C},\;(k_{i},l_{i})\text{ coprime in }\mathbb{Z}^{2},$
where $\Re c_{i}\in[0,1)\,\forall i$, and if $(k_{i},l_{i})=(-k_{j},-l_{j})$ then $c_{i}\neq 1-c_{j}$.
It is easy to see that for every $\eta$ of $\gamma$ type, the space $K\eta$ spanned by it is invariant under $U^{\pm 1}$ and $S^{\pm 1}$, and thus is an $A$-module.
Theorem 2.3 ([Ore30], see also [Sab05, Proposition 1.7]).
The correspondence $\eta\mapsto K\eta$ is an isomorphism between the group of functions of $\gamma$ type and the group of isomorphism classes of 1-dimensional $A$-modules.
Remark 2.4.
The formulations of the theorem in [Ore30] and in [Sab05, Proposition 1.7] are slightly different from ours. For example, in [Sab05, Proposition 1.7] negative powers of $\Gamma(k_{i}u+l_{i}s+c_{i})$ are allowed. On the other hand, [Sab05, Proposition 1.7] reduces the set of allowed pairs $(k_{i},l_{i})$ to include exactly one of $\{(k,l),(-k,-l)\}$ for any coprime $(k,l)\in\mathbb{Z}^{2}$. The equivalence of the two formulations follows from the Euler’s reflection formula:
(12) $\Gamma(x)\Gamma(1-x)=\frac{\pi}{\sin(\pi x)}=\frac{2\pi i}{1-\exp(-2\pi ix)}\exp(-i\pi x)=\phi(x)\exp(-i\pi x),$
where $\phi$ is a periodic function with period 1. Thus multiplication by $\phi(ku+ls+c)\exp(-i\pi c)$ defines an isomorphism between $\langle\Gamma(ku+ls+c)^{-1})\rangle$ and $\langle\Gamma(-ku-ls+1-c)\exp(i\pi ku+i\pi ks)\rangle$.
3. From hypergeometric polynomials to 1-dimensional $A$-modules
Let $P_{n}(z)=\sum c(n,k)z^{k}$ be a family of monic polynomials of rational type. In particular, there exist relatively prime polynomials $N(u,s)$ and $D(u,s)$ such that
(13) $c(n,k+1)D(n,k)\equiv c(n,k)N(n,k)\quad\forall n,k\in\mathbb{Z}$
Recall that we extend the domain of definition of the coefficients $c(n,k)$ to $\mathbb{Z}^{2}$ by setting $c(n,k)=0$ outside of the range $0\leq k\leq n$.
A priori, it is possible that for some $n\geq k\geq 0$ we have $N(n,k)=D(n,k)=0$. Since $D$ and $N$ are co-prime, the set $X=\{(u,s)\in\mathbb{C}^{2}\,|\,N(u,s)=D(u,s)\}$ is finite. For $(n,k)\in\mathbb{Z}^{2}\cap X$, (13) puts no restriction on $c(n,k)$. However, the following lemma implies that (13), together with the condition that the family is quas-orthogonal, uniquely defines the family.
Lemma 3.1 (Appendix A).
Let $\{Q_{n}\}_{n=0}^{\infty}$ be a monic quasi-orthogonal family. Suppose that there exists $N\in\mathbb{Z}_{\geq 0}$ such that $P_{n}=Q_{n}$ for all $n\geq N$. Then $P_{n}=Q_{n}$ for all $n$.
We postpone the proof of this lemma, as well all the lemmas in this section, to Appendix A.
A central role in this paper is played by the rational function
(14) $f(u,s):=\frac{N(u,s)}{D(u,s)}$
Lemma 3.2 (Appendix A).
There exist polynomials $x,y\in\mathbb{C}[u,s]$ such that
(15) $N(u,s)=(u-s)x(u,s),\quad D(u,s)=(s+1)y(u,s).$
By Favard’s theorem (Theorem 2.1 above), there exist (unique) sequences $\{\alpha_{n}\}_{n=1}^{\infty}$ and $\{\beta_{n}\}_{n=1}^{\infty}$ of complex numbers such that $\beta_{n}\neq 0$ for all $n\in\mathbb{Z}_{\geq 1}$ and
(16) $zP_{n}=P_{n+1}+\alpha_{n}P_{n}+\beta_{n}{P_{n-1}}\quad\forall n\in\mathbb{Z}_{\geq 1}$
Lemma 3.3.
There exist rational functions $\alpha$ and $\beta$ such that $\alpha(n)=\alpha_{n}$ and $\beta(n)=\beta_{n}$, except possibly for finitely many $n$.
Let $L$ be the $\mathbb{C}[u,s]$-module of all $\mathbb{C}$-valued functions on $\mathbb{Z}^{2}$, with the action given by restriction of any polynomial in $\mathbb{C}[u,s]$ to $\mathbb{Z}^{2}$, by substitution of $n$ for $u$ and $k$ for $s$. Let $N$ be the extension of scalars $N:=L\otimes_{\mathbb{C}[u,s]}\mathbb{C}(u,s)$.
Recall the notation $K=\mathbb{C}(u,s)$ and $A:=$ the subalgebra of $\operatorname{End}_{\mathbb{C}}(K)$ generated by $K$ and the shift operators $U^{\pm 1}:u\mapsto u\pm 1;\,S^{\pm 1}:s\mapsto s\pm 1$. The $S^{\pm 1},U^{\pm 1}$ act on $K$ and on $L$ by $(U^{\pm 1}g)(u,s):=g(u\pm 1,s)$ and $(S^{\pm 1}g)(u,s):=g(u,s\pm 1)$. Let $S^{\pm 1},U^{\pm 1}$ act on $N$ by $S^{\pm 1}(l\otimes g):=S^{\pm 1}l\otimes S^{\pm 1}g$ and $U^{\pm 1}(l\otimes g):=U^{\pm 1}l\otimes U^{\pm 1}g$. This defines a structure of an $A$-module on $N$. Let $w$ denote the image of the function $c(n,k)$ in $N$ and let $M$ be the $A$-submodule of $N$ generated by $w$.
Lemma 3.4 (Appendix A).
We have $w\neq 0$ and $Sw=fw$.
Theorem 3.5.
$\dim_{K}M=1$.
Proof.
By the previous lemma, $w$ is an eigenvector of $S$. Thus the three-term relation (16) implies that the module $M$ is spanned by $w$ and $U^{-1}w$. Indeed, from $Sw=fw$ we have $S^{-1}w=(S^{-1}f)^{-1}w$. Looking at the coefficient of $z^{k}$ in (16) we have
(17) $(S^{-1}f)^{-1}c(n,k)=(S^{-1}c)(n,k)=Uc(n,k)+\alpha(n)c(n,k)+\beta(n)U^{-1}c(n,k)$
for almost every $n\geq 1$. This also trivially holds for every $n\leq-2$. Thus, if we multiply both sides of the equality by a polynomial in $n$, it will hold for all $n$. Therefore, in the $A$-module $M\subset N=L\otimes_{\mathbb{C}[u,s]}K$ we have
(18) $(S^{-1}f)^{-1}w=S^{-1}w=Uw+\alpha w+\beta U^{-1}w$
Thus $Uw$ is linearly dependent on $w,U^{-1}w$, and thus $w,U^{-1}w$ span $M$, and $\dim_{K}M\leq 2$. Suppose by way of contradiction that $\dim_{K}M=2$. Let $v:=U^{-1}w$. Since $\dim_{K}M=2$, $v$ and $Uv$ span $M$ and are linearly independent. From (18) we have
(19) $U^{2}v=-\beta v+\lambda Uv,$
where $\lambda=S^{-1}f^{-1}-\alpha\in K$. Since $U$ and $S$ commute, $Sv=SU^{-1}w=U^{-1}Sw=U^{-1}(fw)=(U^{-1}f)U^{-1}w=(U^{-1}f)v$. In other words $Sv=\delta v$, where $\delta=U^{-1}f\in K$. Applying $S$ to (19) and using $US=SU$ we have
(20) $SU^{2}v=S(-\beta v)+S(\lambda Uv)=-S(\beta)\delta v+S(\lambda)U(\delta v)=-S(\beta)\delta v+S(\lambda)U(\delta)U(v)$
We also have
(21) $SU^{2}v=U^{2}(Sv)=U^{2}(\delta v)=U^{2}\delta U^{2}v=-U^{2}(\delta)\beta v+U^{2}(\delta)\lambda Uv$
Thus
(22) $-S(\beta)\delta v+S(\lambda)U(\delta)U(v)=-U^{2}(\delta)\beta v+U^{2}(\delta)\lambda Uv$
Since $v,Uv$ are linearly independent, comparing their coefficients we have
(23) $-S(\beta)\delta=-U^{2}(\delta)\beta\text{ and }S(\lambda)U(\delta)=U^{2}(\delta)\lambda$
The first equation implies
(24) $U^{2}(\delta)/\delta=S(\beta)/\beta=1,$
Thus $\delta$ does not depend on $u$, and since $\delta=U^{-1}f$ neither does $f$. However, this contradicts Lemma 3.2, that says that $f$ has the term $u-s$. ∎
Corollary 3.6.
There exists a (unique) rational function $h\in\mathbb{C}(u,s)$ such that outside of the zero set of a polynomial $r(n,k)$ we have
(25) $c(n+1,k)=h(n,k)c(n,k).$
Proof.
Since $\dim_{K}M=1$, there exists a (unique) rational function $h\in\mathbb{C}(u,s)$ such that $Sw=hw.$ Then we have
$fShw=SUw=USw=hUfw,$
thus $fSh=hUf$. By construction of $M$ this implies that (25) holds outside of the zero set of a polynomial $r(n,k)$. ∎
4. Classification of pointed 1-dimensional $A$-modules satisfying the 3-term relation
Let $\Phi(u,s)$ be a meromorphic function of the form $g\phi$, where $\phi$ is of $\gamma$ type and $g\in\mathbb{C}(u,s)$. Then $S\Phi=f\Phi$ for some $f\in\mathbb{C}(u,s)$. Suppose also that $\Phi$ satisfies the 3-term recurrence relation (18):
(26) $S^{-1}\Phi=U\Phi+\alpha\Phi+\beta U^{-1}\Phi,\text{ where }\alpha,\beta\in\mathbb{C}(u)\text{ with }\beta\not\equiv 0.$
By Theorem 3.5, $\Phi$ generates a one-dimensional $A$-module. Thus $U\Phi=h\Phi$ for some $h\in\mathbb{C}(u,s)$. This $h$ necessarily satisfies
(27) $Uf/f=Sh/h$
Since $S\Phi=f\Phi$, we have $S^{-1}\Phi=S^{-1}(f)^{-1}\Phi$, and (26) is equivalent to
(28) $S^{-1}(f)^{-1}=h+\alpha+\beta/U^{-1}(h),$
We are looking for the space of joint solutions of (27) and (28).
Theorem 4.1.
Suppose that $U\Phi=h\Phi,\,\,S\Phi=f\Phi$, and $\Phi\neq 0$ satisfies (26). Then there exist polynomials $g\in\mathbb{C}[u,s]$ and $p,q,w\in\mathbb{C}[t]$, and a rational function $\sigma\in\mathbb{C}(u)$ such that
(29) $f=\frac{q(s+u)p(s-u)Sg}{w(s)g},\quad h=\frac{\sigma(u)q(s+u)Ug}{p(s-u-1)g}.$
Moreover, we can choose $p,q,w,$ and $g$ such that $p,q$ are monic, $g$ is coprime to each of the polynomials $w(s{{-1}}),q(u+s),p(s-u)$ and has no factors independent of $s$. The tuple $(p,q,w,\sigma,g)$ satisfying these conditions is unique, except that $g$ can be multiplied by a constant.
To prove the theorem we will use the following proposition.
Proposition 4.2.
Let $a$ be an irreducible monic non-constant polynomial in $s$ with coefficients in $\mathbb{C}(u)$. If $S^{k}U^{l}a=a$ and $(k,l)\neq(0,0)$ then $l\neq 0$ and $a=s-(k/l)u+c$ with $c\in\mathbb{C}$.
Proof.
First of all $l\neq 0$ since otherwise $S^{k}a=a$, thus $a\in\mathbb{C}(u)$, and since $a$ is monic we have $a=1$, contradicting the assumption that $a$ is not constant. Let $d:=gcd(k,l),k^{\prime}:=k/d,l^{\prime}:=l/d.$ Let $m,n$ be such that $mk^{\prime}+nl^{\prime}=1$. Now, perform a linear change of variables: $x:=l^{\prime}s-k^{\prime}u$, $y:=ms+nu$. Under this change, the shift operator in $y$ becomes $S^{k^{\prime}}U^{l^{\prime}}$.
The condition $S^{k}U^{l}a=a$ implies that $a$ is invariant to shifts $y\mapsto y+d$. Thus $a$ depends only on $x$. Thus $a$ is a polynomial in the expression $ls-ku$ with complex coefficients. Since it is irreducible it must be linear. Since it is monic, it must be $s-(k/l)u+c$. ∎
Proof of Theorem 4.1.
Since $\Phi=g\phi$, where $\phi$ is of $\gamma$ type and $g\in\mathbb{C}(u,s)$, $\Phi$ has the form
$\Phi=\exp(cu+ds)g(s,u)\prod\Gamma(k_{i}u+l_{i}s+c_{i}),$
with $k_{i},l_{i}\in\mathbb{Z}$ and $(k_{i},l_{i})$ coprime for every $i$, $\Re c_{i}\in[0,1)$ and if $(k_{i},l_{i})=(-k_{j},-l_{j})$ then $c_{i}\neq 1-c_{j}$. By removing the condition $\Re c_{i}\in[0,1)$ and using the identity $x\Gamma(x)=\Gamma(x+1)$ we assume without loss of generality that for every $i$, $k_{i}u+l_{i}s+c_{i}$ is not a factor in the numerator of $g$, and $k_{i}u+l_{i}s+c_{i}-1$ is not a factor in the denominator of $g$.
We claim that for every $i$ there are only five possibilities for the pair $(k_{i},l_{i})$. Namely, $(k_{i},l_{i})\in\{(\pm 1,1),(0,-1),(\pm 1,0)\}$ for every $i$. Indeed, fix $i$. If $l_{i}=0$ then $k_{i}\in\{\pm 1\}$ since $(k_{i},l_{i})$ are coprime. Thus we assume $l_{i}\neq 0$. Assume also, without loss of generality, that for every $j$ with $(k_{j},l_{j})=(l_{i},l_{i})$, we have $\Re c_{j}\leq\Re c_{i}$. Consider all $n\in\mathbb{Z}$ such that $\Phi$ has a pole along the line $k_{i}u+l_{i}s+c_{i}=n$. The maximal $n$ will be 00. For $S^{-1}\Phi$, $U\Phi$, and $U^{-1}\Phi$ such maximal $n$’s will be $l_{i},-k_{i}$, and $k_{i}$ respectively. Every pole in (26) should occur in at least two terms. The terms $S^{-1}\Phi,U\Phi$, and $U^{-1}\Phi$ enter with non-zero coefficients that cannot cancel the pole, and one of the numbers $k_{i},-k_{i}$ is non-negative. Thus the two largest of the numbers $(0,l_{i},k_{i},-k_{i})$ are equal. Thus either $k_{i}=0$ and $l_{i}<0$, or $l_{i}=|k_{i}|$. Since $k_{i}$ and $l_{i}$ are coprime, this implies that $(k_{i},l_{i})\in\{(0,-1),(\pm 1,1)\}$.
The pairs (±$(\pm 1,1)$ give the factors $s\pm u+c_{i}$ in $f=S\Phi/\Phi$, the pairs (±$(\pm 1,0)$ have no effect on $f$, and the pair $(0,-1)$ produces the factor $(s-1+c_{i})^{-1}$ in $f$. The term $\exp(cu+ds)$ produces a multiplicative constant $d$. The assumption that $k_{i}u+l_{i}s+c_{i}-1$ is not a term in the numerator of $g$ for any $i$ implies that the numerator of $g$ has no common factors with $w(s{{-1}}),q(u+s),p(s-u)$. Thus $f$ has the form as in (29), and by (27) so does $h$, except that $g$ is allowed to be a rational function.
It is left to prove that $g$ is a polynomial. We can write
(30) $\sigma=\delta_{1}/\gamma,\quad\alpha=\delta_{0}/\gamma,\quad\beta/U^{-1}\sigma=\delta_{-1}/\gamma,$
where $\gamma,\delta_{1},\delta_{0},\delta_{-1}\in\mathbb{C}[u]$ are polynomials that have no overall common factor. By (29), equation (28) becomes
(31) $\gamma(u)w(s-1)S^{-1}(g)=\delta_{1}(u)q(s+u)q(s+u-1)Ug\,+\\ +\delta_{0}(u)q(s+u-1)p(s-u-1)g+\delta_{-1}(u)p(s-u)p(s-u-1)U^{-1}g.$
Suppose, by way of contradiction, that $g$ is not a polynomial, and let $a$ be an irreducible factor of the denominator of (the reduced expression for) $g$. By our assumption on $g$, $a$ depends on $s$. Then $S^{-1}a$ is an irreducible factor in the denominator of $S^{-1}g$. The assumption that $k_{i}u+l_{i}s+c_{i}-1$ is not a factor in the denominator of $g$ for every $i$ implies that the denominator of $g$ is coprime to $w(s)$. Thus $w(s-1)$ cannot cancel the factor $S^{-1}a$, and thus $S^{-1}a$ appears in the denominator of the reduced expression for the LHS of (31). Thus $S^{-1}a$ must appear in the denominator of the reduced expression for one of the terms of the RHS as well. Therefore $S^{-1}a=U^{i}b$, where $i\in\{-1,0,1\}$ and $b$ is another factor in the denominator of $g$. Then $S^{-1}b$ is also a factor in the denominator of $S^{-1}g$. Since the number of factors is finite, the factors obtained in this way must repeat. It follows that $S^{-k}U^{l}a=a$, where $k,l\in\mathbb{Z}$, $k\geq|l|$, and $k>0$.
Thus, by Proposition 4.2, $a=d(u+(l/k)s+c)$ for some $c,d\in\mathbb{C}$. Without loss of generality we assume $d=1$ and that among all the factors of the form $u+(l/k)s+e$ with the same ratio $l/k$, $a$ has minimal $\Re e$. The third term of the RHS of (31) has $U^{-1}a=a-1$ in the denominator. Thus $a-1\in\{S^{-1}b^{\prime},b^{\prime},Ub^{\prime}\}$ for some factor $b^{\prime}$ of the denominator of $g$. But then $b\in\{a-1+l/k,a-1,a-2\}$. From the assumption on $a$ and the inequality $k\geq|l|$, we get $l/k=1$ and thus $a=u+s+c$.
We proved that the denominator of $g$ can only have factors of the form $u+s+c$. Let $a^{\prime}$ be the factor with maximal $\Re c$. Then the first term of the RHS has $Ua^{\prime}=a^{\prime}+1$ in the denominator, and no other term of (31) has this factor in the denominator - contradiction.
For the uniqueness statement, suppose we have 2 tuples $(p_{i},q_{i},w_{i},g_{i}),\,i\in\{1,2\}$ such that for each $i$, $p_{i},q_{i}$ are monic, $g_{i}(s,u)$ is coprime to $p_{i}(s-u)$, to $q_{i}(s+u)$ and to $w_{i}(s)$ and has no factors independent of $s$, and
(32) $\frac{q_{1}(s+u)p_{1}(s-u)Sg_{1}}{w_{1}(s)g_{1}}=\frac{q_{2}(s+u)p_{2}(s-u)Sg_{2}}{w_{2}(s)g_{2}}$
We claim that $p_{1}=p_{2}$, $q_{1}=q_{2}$, $w_{1}=w_{2}$, and $g_{1}/g_{2}$ is a constant. We prove this by induction on $d:=\min(\deg g_{1},\deg g_{2})$. The base case $d=0$ is straightforward. For the induction step, assume WLOG $\deg g_{1}\geq\deg g_{2}$ and let $a$ be an irreducible factor in $g_{1}$. Since $g$ is coprime to $p(s-u),q(s+u)$, $a$ appears in the denominator of the reduced expression for the LHS, and thus also of the reduced expression for the RHS. Thus $a$ divides either $g_{2}(s,u)$ or $w(s)$. If $a$ divides $g_{2}$ then we can cancel it out from both $g-1$ and $g_{2}$, and the statement holds by induction. Assume now that $a$ divides $w(s)$. Then $a$ depends only on $s$, and so does $Sa$, which is a factor in the numerator of the reduced expression, since $Sg_{1}$ is coprime to $w_{1}(s)$. Thus $Sa$ divides $Sg_{2}$ and thus $a$ divides $g_{2}$, and the statement again follows by induction. ∎
We will show below that there are only finitely many possibilities for degrees of $p,q,w$, in particular we will prove that these degrees are at most $2$. The key tool is the following lemma.
Lemma 4.3.
In the setting of Theorem 4.1, define $d_{1}:=\deg p,\;d_{2}:=\deg q,\;d_{3}:=\deg w$; then the four polynomials $a_{1}(x)=(x-1)^{2d_{1}},\;a_{2}(x)=(x-1)^{d_{1}}(x+1)^{d_{2}},\;a_{3}(x)=(x+1)^{2d_{2}},\;a_{4}(x)=x^{d_{3}}$ are linearly dependent over $\mathbb{C}$.
Proof.
Let us first explain the main idea. Let $\mathcal{R}[u]$ be the polynomial ring over a commutative $\mathbb{C}$-algebra $\mathcal{R}$, and consider the (discontinuous) map $\mathcal{R}[u]\to\mathcal{R}$ which assigns to each polynomial its leading coefficient. It is easy to see if the polynomials ${A_{1}},\ldots,{A_{n}}$ are dependent over $\mathbb{C}$ then their leading coefficients $a_{1},\ldots,a_{n}$ are also dependent over $\mathbb{C}$.
Indeed, suppose have a non-trivial relation $\sum_{i=1}^{n}c_{i}A_{i}=0$. Without loss of generality we may assume that we have $c_{i}\neq 0$ for all $i$, that the polynomials $A_{1},\ldots,A_{k}$ have degree $d$, and that $A_{i},i>k$ have strictly smaller degree. Then by examining the coefficient of $u^{d}$, we see that the relation $\sum_{i=1}^{n}c_{i}A_{i}=0$ implies the relation $\sum_{i=1}^{k}c_{i}a_{i}=0$.
For our application we need a slight generalization. Suppose that $\mathcal{R}$ is also an integral domain, that $B_{1},\ldots,B_{n}$ are polynomials with the same leading coefficient $b$, and that $A_{1}B_{1},\ldots A_{n}B_{n}$ are linearly dependent, then we claim that $a_{1},\ldots,a_{n}$ are dependent over $\mathbb{C}$ as before. Indeed, the leading coefficient of $A_{i}B_{i}$ is $a_{i}b$, and arguing as above we obtain a linear relation $\sum_{i=1}^{k}a_{i}b=0$. Since $\mathcal{R}$ is a domain we can simply cancel $b$ from this relation.
Now for the proof of the lemma we substitute $s=xu$ in (31) to obtain the identity
(33) $A_{1}g(xu,u+1)+A_{2}g(xu,u)+A_{3}g(xu,u-1)+A_{4}g(xu-1,u)=0$
where
$A_{1}=\delta_{-1}(u)p(xu-u)p(xu-u-1),\quad$ $A_{2}=\delta_{0}(u)p(xu-u-1)q(xu+u-1),$
$A_{3}=\delta_{1}(u)q(xu+u)q(xu+u-1),\quad$ $A_{4}=-\gamma(u)w(xu-1).$
It is easily seen that for each $i$, the leading $u$-coefficient of $A_{i}$ is a non-zero $\mathbb{C}$-multiple of $a_{i}(x)$, while each of the four $g$-factors (33) has the same leading $u$-coefficient. Thus the result follows. ∎
In other words there exist constants $c_{1},c_{2},c_{3},c_{4}\in\mathbb{C}$, not all 00, such that
(34) $c_{1}(x-1)^{2d_{1}}+c_{2}(x-1)^{d_{1}}(x+1)^{d_{2}}+c_{3}(x+1)^{2d_{2}}+c_{4}x^{d_{3}}\equiv 0.$
Lemma 4.4.
We have $|d_{1}-d_{2}|\leq 1$.
Proof.
Replacing $k$ by $-k$ in (34) we can assume $d_{1}\geq d_{2}$. Thus we have to show $d_{1}\leq d_{2}+1$.
Suppose by way of contradiction that $d_{1}\geq d_{2}+2$. Then the polynomial $(k-1)^{2d_{1}}$ has, among others, monomials of degrees $2d_{1}$ and $12d_{1}-$ 1. No term of $(k+1)^{2d_{2}}$ or of $(k-1)^{d_{1}}(k+1)^{d_{2}}$ can cancel any of these, and the term $k^{d_{3}}$ can only cancel one of them. Thus we must have $c_{1}=0$. But then, the polynomial $(k-1)^{d_{1}}(k+1)^{d_{2}}$ has terms of degrees $d_{1}+d_{2}$ and $d_{1}+d_{2}-1$ and again $k^{d_{3}}$ can only cancel one of them. Thus we have $c_{1}=c_{2}=0$ and thus $c_{3},c_{4}\neq 0$. If $d_{2}>0$ then $(k+1)^{2d_{2}}$ has at least three terms, and they cannot be cancelled by $c_{4}k^{d_{3}}$. Thus we must have $d_{2}=d_{3}=0$. Then looking at the degree of $s$ in (31), we see that the term $\delta_{-1}p(s-u)p(s-u-1)U^{-1}g$ has higher degree in $s$ then the other terms. This is a contradiction. ∎
Lemma 4.5.
Suppose that $d_{1},d_{2}\geq 1$. Then $d_{1}=d_{2}\in\{1,2\}$.
Proof.
Suppose $c_{4}=0$ in (34). Then setting $k=\pm 1$ we deduce $c_{1},c_{3}=0$ and hence $c_{2}=0$ also. Thus we may assume $c_{4}\neq 0$. Now setting $k=\pm 1$ we get $c_{1},c_{3}\neq 0$.
Let $\lambda$ be a root of the quadratic polynomial $c_{1}x^{2}+c_{2}x+c_{3}$, and let $\mu$ be a root of the polynomial $(k-1)^{d_{1}}-\lambda(k+1)^{d_{2}}$. Then setting $k=\mu$ in (34) the sum of the first three terms vanish and we get $c_{4}\mu^{d_{3}}=0$, which implies $\mu=0$ (and $d_{3}\geq 1$). Thus 0 is the only root of $(k+1)^{d_{1}}-\lambda(k-1)^{d_{2}}$, which means this polynomial is a monomial:
(35) $(k-1)^{d_{1}}-\lambda(k+1)^{d_{2}}=\varepsilon k^{d_{4}}\quad\text{ for some }\varepsilon,d_{4}$
If $d_{4}=0$ then $d_{1}=d_{2}=1$, so we may assume $d_{4}>0$. Setting $k=0,1,-1$ in (35) we get
$\lambda=(-1)^{d_{1}},\,\varepsilon=(-1)^{d_{1}+1}2^{d_{2}},\,\varepsilon=(-1)^{d_{1}+d_{4}}2^{d_{1}}.$
This forces $d_{1}=d_{2}=:d$, and (35) becomes
(36) $(k-1)^{d}-(-1)^{d}(k+1)^{d}=(-1)^{d+d_{4}}2^{d}k^{d_{4}}$
The coefficient of $k$ on the left is $2d(-1)^{d-1}$ which implies $d_{4}=1$ and $2d=2^{d}$. Thus $d\in\{1,2\}$. ∎
Lemma 4.6.
If $d_{1}=d_{2}=2$ then $d_{3}=2$ and $\delta_{1}+\delta_{0}+\delta_{-1}=0$.
Proof.
Substituting $d_{1}=d_{2}=2$ in (34) we obtain
(37) $-c_{4}k^{d_{3}}=(c_{1}+c_{2}+c_{3})k^{4}+4(c_{3}-c_{1})k^{3}+2(3c_{1}-c_{2}+3c_{3})k^{2}+4(c_{3}-c_{1})k+(c_{1}+c_{2}+c_{3})$
Since the LHS is a monomial, so is the RHS. Since the coefficients $4(c_{3}-c_{1})$ and $c_{1}+c_{2}+c_{3}$ appear twice, both must vanish. Thus, $c_{3}=c_{1}$, $c_{2}=-2c_{1}$, and $d_{3}=2$. Thus the top degree terms in $s$ of the RHS of (31) must cancel, and thus $\delta_{1}+\delta_{0}+\delta_{-1}=0$. ∎
Summarizing the last three lemmas we obtain the following corollary.
Corollary 4.7.
In the setting of Theorem 4.1, denote $d_{1}:=\deg p,\,d_{2}:=\deg q,d_{3}:=\deg w$.
Then we have $d_{3}\leq 2$, and either $d_{1},d_{2}\leq 1$ or $d_{1}=d_{2}=d_{3}=2$.
Proof.
If $d_{1},d_{2}\leq 1$ then the degree in $s$ of the RHS of (31) is at most $\deg g+2$. Thus $d_{3}\leq 2$. Suppose now $\max(d_{1},d_{2})\geq 2$. Then, by Lemma 4.4, $\min(d_{1},d_{2})\geq 1$, and Lemma 4.5 implies $d_{1}=d_{2}=2$. By Lemma 4.6 this implies $d_{3}\leq 2$.
Altogether, we obtain $d_{3}\leq 2$, and either $d_{1},d_{2}\leq 1$, or $d_{1}=d_{2}=2$. ∎
In the special case $g=1$ we can further restrict the options for $p,q,$ and $w$.
Lemma 4.8.
If $d_{1}=d_{2}=2$, $w(-1)=0$, and $w$ is monic then one of the following holds.
  1. (a)
    $w(s)=(s+1/2)(s+1)$, $p(t)=q(-1-t)$
  2. (b)
    $w(s)=(s+1)(s+3/2)$, $p(t)=q(-2-t)$
Proof.
Shifting $u$ and $s$ we can assume $p(t)=t^{2}+c$, $q(t)=t^{2}+d$ for some $c,d$. By Lemma 4.6 we have $d_{3}=2$ and $\delta_{1}+\delta_{0}+\delta_{-1}=0$. Since $\deg w=2$, the term of $s^{3}$ in the RHS of (31) must vanish. But this term is $4(\delta_{1}-\delta_{-1})u-2(\delta_{1}+2\delta_{0}+\delta_{-1})$. Since $\delta_{1}+\delta_{0}+\delta_{-1}=0$, we obtain that $\delta_{0}=2(\delta_{1}-\delta_{-1})u$, and thus $\delta_{1}(2u+1)=\delta_{-1}(2u-1)$. Thus we can assume without loss of generality that $\delta_{-1}=2u+1$ and $\delta_{1}=2u-1$. This implies $\delta_{0}=-4u$.
Then the coefficient of $s^{2}$ in the RHS of (31) is $32u^{3}-8u+2(d-c)$. This does not vanish identically, thus $w$ must have degree 2. Then we can assume $w$ is monic and $\gamma(u)=32u^{3}-8u+2(c-d)$.
The coefficient of $s$ in the RHS of (31) is $2(d-c)+12u-16u^{2}(c-d)-48u^{3}$. Since this has to be divisible by $\gamma$, we have $c=d$. In this case $w(s-1)=(s-1)(s-1/2)$, $\gamma=8u(4u^{2}-1)$. Thus $w(s)=s(s+1/2)$.
Shifting back in $s$ we get the same shift for $p$ and $q$, while shifting in $u$ we obtain opposite shifts. Since $w(-1)=0$ after the shift, we can shift $s$ either by $1/2$ or by $1$. If we shift by $1/2$ we get case (a), and if we shift it by $1$ we get case (b). ∎
This lemma, together with Corollary 4.7, implies the following corollary.
Corollary 4.9.
In the setting of Theorem 4.1, suppose that $g=1$, $w(-1)=0$, and $w$ is monic. Then one of the following holds.
  1. (a)
    $\deg p,\deg q\leq 1$
  2. (b)
    $w(s)=(s+1/2)(s+1)$, $p(t)=q(-1-t)$
  3. (c)
    $w(s)=(s+1)(s+3/2)$, $p(t)=q(-2-t)$
5. Proof of Theorems A and B
Lemma 5.1 (Direct computation).
Let
(38) $Q_{u}(z):=\frac{(-1)^{n}\prod_{k=1}^{j}(\delta_{k})_{(n)}}{\prod_{l=1}^{i}(\gamma_{l})_{(n)}}\cdot{\,}_{i}F_{j}\left(\begin{matrix}-u&&\gamma\\ &\delta\end{matrix};z\right)$
Then $SQ_{u}=fQ_{u}$, where
(39) $f(u,s)=\frac{s-u}{s+1}\frac{\prod_{l=1}^{i}(s+\gamma_{l})}{\prod_{k=1}^{j}(s+\delta_{k})}$
Conversely, if $f\in\mathbb{C}(u,s)$ is a rational function given by the formula (39), and $R_{n}(z)$ is a monic polynomial family of rational type given by $f$, then $R_{n}(z)=Q_{n}(z)$.
The 5 hypergeometric polynomial families in Theorem A are $Q_{n}(z)$, with the parameters $\delta$ and $\gamma$ given by the following table.
N Name $\gamma$ $\delta$
1 Jacobi $n+a$ $b$
2 Laguerre $\emptyset$ $b$
3 Bessel $n+a$emptyset
4 $E_{n}^{(c)}$ $(1-n-c,n+1,n+c)$ $1/2$
5 $F_{n}^{(c)}$ $(1-n-c,n+2,n+c+1)$ $3/2$
The area of definition is $a\notin\mathbb{Z}_{<0}$; $b\notin\mathbb{Z}_{\leq 0}$; $c\notin\mathbb{Z}_{<0}$, and in addition for $E_{n}$: $c\neq 0$ and for $j_{n}^{b-1,a-b}$: $b-a\notin\mathbb{Z}_{>0}$.
For the proof of Theorem B we will need one more lemma, that is also straightforward.
Lemma 5.2.
Let $Q_{n}$ be a polynomial family satisfying $c_{Q}(n+1,k)/c_{Q}(n,k)=f_{Q}(n,k)$ for some $f_{Q}\in\mathbb{C}(u,s)$. Let $g\in\mathbb{C}[u,s]$ be a polynomial and define a new family $R_{n}$ by $c_{R}(n,k):=g(n,k)c_{Q}(n,k)$. Then $R_{n}=g(n,z\partial_{z})Q_{n}$, and $c_{R}(n+1,k)/c_{R}(n,k)=f_{Q}(n,k)g(n,k+1)/g(n,k)$.
5.1. Proof of Theorem B
By Lemmas 5.1 and 5.2, Theorem B is equivalent to the following one.
Theorem 5.3.
Let $\mathcal{P}$ be a monic polynomial family of rational type, and let $f\in\mathbb{C}(u,s)$ be such that $S\mathcal{P}=f\mathcal{P}$. Then there a polynomial $g\in\mathbb{C}[u,s]$ such that the reduced form of $f$ is
(41) $f=\frac{Sg}{g}\cdot\frac{s-u}{s+1}\cdot\frac{p_{1}(s-u)q(u+s)}{w_{1}(s)},$
where $p_{1},q,w_{1}\in\mathbb{C}[t]$ are polynomials in one variable that belong to the following list
  1. (a)
    $w_{1}(s)=s+b,\quad p_{1}(t)=1,\quad q(t)=t+a$ for some $a,b\in\mathbb{C}$
  2. (b)
    $w_{1}(s)=s+b,\quad p_{1}(t)=1,\quad q(t)=1$ for some $b\in\mathbb{C}$
  3. (c)
    $w_{1}(s)=1,\quad p_{1}(t)=1,\quad q(t)=t+a$ for some $a\in\mathbb{C}$
  4. (d)
    $w_{1}(s)=s+b,\quad p_{1}(t)=t+d,\quad q(t)=(t+a)(t+c)$ for some $a,b,c,d\in\mathbb{C}$
Proof.
Let $M$ be the $A$-module defined by $\mathcal{P}$ in §3. By Theorem 3.5, it is one-dimensional. By abuse of notation, we will denote the element of $M$ defined by $\mathcal{P}$ also by $\mathcal{P}$. Let $f,h\in\mathbb{C}(u,s)$ be such that $S\mathcal{P}=f\mathcal{P}$ and $U\mathcal{P}=h\mathcal{P}$. By Theorems 2.3 and 4.1 we have
(42) $f=\frac{q(u+s)p(s-u)Sg}{w(s)g}$
where $g$ is coprime to $p(s-u),q(s+u)$ and $w(s-1)$.
The polynomial family $\mathcal{P}$ is uniquely defined by $f$, which in turn, by (42), is defined by $p,q$, and $w$. By Lemma 3.2 we have $p(0)=0$ and $w(-1)=0$. Thus $\deg p,\deg w\geq 1$ and $p=(u-s)p_{1}$, $w=(s+1)w_{1}$ for some polynomials $p_{1},w_{1}$. By Corollary 4.7 we have $\deg p,\deg w\leq 2$. Using the rescaling $z\mapsto cz,f\mapsto cf$ discussed in the introduction, we can assume that $p,q,$ and $w$ are monic.
Consider first the case $\deg p=2$. By Corollary 4.7 in this case we have $\deg q=\deg w=2$. Then $\deg w_{1}=\deg p_{1}=1$ and thus $w_{1}=s+b,\,p_{1}=t+d$ for some $b,d\in\mathbb{C}$. Since $\deg q=2$, we have $q(t)=(t+a)(t+c)$ for some $a,c\in\mathbb{C}$. This is case (d).
It is left to consider the case $\deg p=1$, i.e. $p_{1}=1$. By Corollary 4.7, in this case $\deg q\leq 1$, $\deg w_{1}\leq 1$.
Case 1. $\deg q=1$.
Since $q$ is monic, we have $q(t)=t+a$. Since $\deg w_{1}\leq 1$, we have either $w_{1}(s)=s+b$ or $w_{1}(s)=1$ for some $b\in\mathbb{C}$. These are cases (a) and (c).
Case 2. $\deg q=0,\,\,\deg w_{1}=1$.
Then $w_{1}=s+b$, and $q=1$. This is case (b).
Case 3. $\deg q=0,\,\deg w_{1}=0$.
Then $w_{1}=1$, and $q=1$. In this case, the degree in $s$ of the LHS of (31) is $\deg_{s}g+1$, while the degree in $s$ of the RHS is $\deg_{s}g+2$. This is a contradiction, and thus this case is not possible. ∎
5.2. Proof of Theorem A
Let $f$ be the rational function defining $\mathcal{P}$. By Theorem 5.3 there a polynomial $g\in\mathbb{C}[u,s]$ such that the reduced form of $f$ is
(43) $f=\frac{Sg}{g}\cdot\frac{s-u}{s+1}\cdot\frac{p_{1}(s-u)q(u+s)}{w_{1}(s)},$
where $p_{1},q_{1},w_{1}\in\mathbb{C}[t]$ are polynomials in one variable that belong to a very restricted list.
By the assumptions on $f$ in the definition of Jacobi type, $g$ does not depend on $u$. Thus $Ug=U^{-1}g=g$. Let us show that $g$ is constant. Suppose by way of contradiction that $g$ is not constant, and let $c$ be a root of $g$ that is not a root of $S^{-1}g$. Substitute $s:=c$ into the 3-term relation (31). Then the right-hand side vanishes, while the left-hand side does not. This is a contradiction, thus $g$ must be constant, and thus we can assume $g=1$.
In cases (a)-(c) of Theorem 5.3, we get that $\mathcal{P}$ is the Jacobi, Laguerre, or Bessel family respectively.
Consider now the case (d) of Theorem 5.3. Then $p=t(t+d)$ for some $d\in\mathbb{C}$. Define $c:=1-d$. By Corollary 4.9, either $w(s)=(s+1)(s+1/2)$ and $q(t)=(t+1)(t+c)$, or $w(s)=(s+1)(s+3/2)$ and $q(t)=(t+2)(t+c+1)$. Thus $f$ equals either the $f$ of $E_{n}^{(c)}$ or the $f$ of $F_{n}^{(c)}$. Thus $P_{n}$ is either $E_{n}^{(c)}$ or $F_{n}^{(c)}$. $\Box$
Summarizing, we have the following table
family $w(s)$ $p(t)$ $q(t)$
Jacobi $(s+1)(s+b)$ $t$ $t+a$
Laguerre $(s+1)(s+b)$ $t$ $1$
Bessel $s+1$ $t$ $t+a$
$E_{n}^{(c)}$ $(s+1)(s+1/2)$ $t(t+1-c)$ $(t+1)(t+c)$
$F_{n}^{(c)}$ $(s+1)(s+3/2)$ $t(t+1-c)$ $(t+2)(t+c+1)$
6. Further properties and details of families of Jacobi type
Let $\mathcal{P}$ be a family of Jacobi type.
Lemma 6.1 (Appendix A.1).
  1. (i)
    $D(k)\neq 0$ for any $k\in\mathbb{Z}_{\geq 0}$.
  2. (ii)
    For all $n\geq k\geq 0$ we have $c(n,k)\neq 0$, and for all $n>k\geq 0$ we have $N(n,k)\neq 0$.
  3. (iii)
    For every $n\geq k\geq 0$ we have $c(n,k)=\prod_{i=k}^{n-1}f(n,i)^{-1}$.
Lemma 6.2 (Appendix A.1).
The rational function $\alpha$ and $\beta$ have finite values at all integers $n\geq 1$. These values coincide with the numbers $\alpha_{n}$ and $\beta_{n}$ in the three-term recursion (16), with one exception: for the Bessel family with $a=0$, $\beta(1)\neq\beta_{1}$. Furthermore, we have
(45) $\alpha(u)=f(u,u-1)^{-1}-f(u+1,u)^{-1}$
and
(46) $\beta(u)=f(u,u-2)^{-1}f(u,u-1)^{-1}-f(u+1,u-1)^{-1}f(u+1,u)^{-1}-\\ -f(u,u-1)^{-2}+f(u,u-1)^{-1}f(u+1,u)^{-1}$
From Lemma 6.2, equation (43) with $g=1$ and equation (44) we obtain the following table for $\alpha$ and $\beta$ for each family of Jacobi type.
family $\alpha$ $\beta$
Jacobi $\frac{2u^{2}+2au+b(a-1)}{(2u+a-1)(2u+a+1)}$ $\frac{u(u+a-1)(u+b-1)(u+a-b)}{(2u+a)(2u+a-1)^{2}(2u+a-2)}$
Laguerre $2u+b$ $u(u+b-1)$
Bessel $\frac{a-1}{(2u+a-1)(2u+a+1)}$ ${{-}}\frac{u(u+a-1)}{(2u+a)(2u+a-1)^{2}(2u+a-2)}$
$E_{n}^{(c)}$ $-\frac{1}{2(2u+c-1)(2u+c+1)}$ $\frac{1}{16(2u+c)(2u+c-1)^{2}(2u+c-2)}$
$F_{n}^{(c)}$ $-\frac{1}{2(2u+c)(2u+c+2)}$ $\frac{1}{16(2u+c+1)(2u+c)^{2}(2u+c-1)}$
Each family is determined by the sequences $\alpha_{n}$ and $\beta_{n}$ and by the scalar $c_{0}=P_{1}(0)$. For all the families except $E_{n}$ we have $c_{0}=-\alpha(0)$. Equivalently, for these families instead the initial condition $P_{1}=z+c_{0}$ we could have used the condition $P_{-1}=0$, and extend the function $\alpha$ to 00. However, for the family $E_{n}^{(c)}$ we have $c_{0}=1/(4c(c+1))\neq 1/(2(c-1)(c+1))=-\alpha(0)$. If we set $c_{0}$ to be $-\alpha(0)$, keeping the functions $\alpha$ and $\beta$ of $E_{n}^{(c)}$ we obtain the family $F_{n}^{(c-1)}$.
7. Investigation of the two new families $E_{n}^{(c)},F_{n}^{(c)}$
7.1. Preliminaries on Lommel polynomials
Let us give some preliminaries on Lommel polynomials from [Wat44, §9] and [Ism09, §6.5]. They are defined by the recursive relation
(48) $2z(n+c)h_{n}^{(c)}=h_{n+1}^{(c)}(z)+h^{(c)}_{n-1}(z),\quad\quad h_{-1}^{(c)}(z)\equiv 0,\,\,h^{(c)}_{0}\equiv 1$
Explicitly, we have [^1] M. Ismail kindly informed us of a typo in [Ism09, (6.5.8)]: $z/2$ should be $2z$.
(49) $h_{2n}^{(c)}(z)=(-1)^{n}\sum_{k=0}^{n}\binom{n+k}{2k}(n+c-k)_{(2k)}(-2z)^{2k},\\ h_{2n+1}^{(c)}(z)=(-1)^{n}\sum_{k=0}^{n}\binom{n+k+1}{2k+1}(n+c-k)_{(2k+1)}(-2z)^{2k+1}$
To describe the discrete measure on the real line for which these polynomials are orthogonal we will need some notation. Let $J_{v}(z)$ denote the modified Bessel function [Wat44]:
(50) $J_{v}(z)=\sum_{n=0}^{\infty}\frac{(-1)^{n}(z/2)^{c+2n}}{\Gamma(n+\nu+1)n!}$
Let $\{j_{\nu,k}\}_{k=1}^{\infty}$ denote the increasing sequence of positive zeroes of $J_{\nu,k}$ on $\mathbb{R}$. Denote $a_{\nu,k}:=j_{\nu,k}^{-1}$.
Theorem 7.1 ([Ism09, (6.5.17)]).
For every $c\in\mathbb{R}_{>0}$ and every $n,m\in\mathbb{Z}_{\geq 0}$ we have
(51) $\sum_{k=1}^{\infty}a_{c-1,k}^{2}(h_{n,c}(a_{c-1,k})h_{m,c}(a_{c-1,k})+h_{n,c}(-a_{c-1,k})h_{m,c}(-a_{c-1,k}))=\frac{\delta_{m,n}}{2(n+c)}$
7.2. Discrete measure for which $E_{n}^{(c)}(z)$ and $F_{n}^{(c)}(z)$ form orthogonal families
Spelling out the definition of $E_{n}$ and $F_{n}$ as hypergeometric functions, we have
(52) $E_{n}^{(c)}(z)=\frac{1}{2^{2n}c_{(2n)}}\sum_{k=0}^{n}\binom{n+k}{2k}(n+c-k)_{(2k)}2^{2k}z^{k},\\ F_{n}^{(c)}(z)=\frac{1}{2^{2n}c_{(2n+1)}}\sum_{k=0}^{n}\binom{n+k+1}{2k+1}(n+c-k)_{(2k+1)}2^{2k}z^{k}$
Combining (52) and (49) we have for any $n\in\mathbb{Z}_{\geq 0}$:
(53) $E_{n}^{(c)}(-z^{2})=\frac{(-1)^{n}}{2^{2n}c_{(2n)}}h_{2n}^{(c)}(z),\quad zF_{n}^{(c)}(-z^{2})=\frac{(-1)^{n}}{2^{2n+1}c_{(2n+1)}}h_{2n+1}^{(c)}(z).$
Theorem 7.1 and (53) imply the following corollary.
Corollary 7.2.
For every $c\in\mathbb{R}_{>0}$ and every $n,m\in\mathbb{Z}_{\geq 0}$ we have
(54) $4c\sum_{k=1}^{\infty}a_{c-1,k}^{2}E_{n}^{(c)}(-a^{2}_{c-1,k})E_{m}^{(c)}(-a^{2}_{c-1,k})=\frac{c}{2^{4n}c^{2}_{(2n)}(2n+c)}\delta_{m,n}$
and
(55) $16c^{2}(c+1)\sum_{k=1}^{\infty}a_{c-1,k}^{4}F_{n}^{(c)}(-a^{2}_{c-1,k})F_{m}^{(c)}(-a^{2}_{c-1,k})=\frac{c+1}{2^{4n}(c+1)^{2}_{(2n)}(2n+1+c)}\delta_{m,n}$
The arguments of [Ism09, §6.5] show that (55) continues to hold for every non-zero $c>-1$.
7.3. Differential equations
Being hypergeometric functions of type $\,{}_{4}F_{1}$, the $E_{n}^{(c)}$ and $F_{n}^{(c)}$ satisfy the following 4th order differential equations.
(56) $(D(D-1/2)-z(D-n)(D-n-c+1)(D+n+c)(D+n+1))E_{n}^{(c)}=0$
and
(57) $(D(D+1/2)-z(D-n)(D-n-c+1)(D+n+c+1)(D+n+2))F_{n}^{(c)}=0,$
where $D=z\frac{\partial}{\partial_{z}}$ is the Euler operator.
Appendix A Proof of technical lemmas
Recall that $\alpha_{n}$ and $\beta_{n}$ denote the coefficients in the 3-term recursion for $P_{n}$, i.e. for every $n>0$ we have
(58) $zP_{n}=P_{n+1}+\alpha_{n}P_{n}+\beta_{n}{P_{n-1}}$
Proof of Lemma 3.1.
Let $\gamma_{n}$, $\mu_{n}$ be sequences such that
(59) $zQ_{n}=Q_{n+1}+\gamma_{n}Q_{n}+\mu_{n}{Q_{n-1}}.$
Let us prove by descending induction on $n$ that $P_{n}=Q_{n}$. The base is $n\in\{N,N+1\}$. For the induction step assume $P_{n}=Q_{n}$ and $P_{n+1}=Q_{n+1}$. Then from (58,59) we have
(60) $zP_{n}-\alpha_{n}P_{n}-\beta_{n}{P_{n-1}}=P_{n+1}=Q_{n+1}=zQ_{n}-\gamma_{n}Q_{n}-\mu_{n}{Q_{n-1}}$
Since $P_{n}=Q_{n}$ this implies
(61) $(\alpha_{n}-\gamma_{n})P_{n}+\beta_{n}{P_{n-1}}-\mu_{n}{Q_{n-1}}=0$
Since $P_{n}$ has degree $n$ and $P_{n-1},Q_{n-1}$ have degree $n-1$, we have $\alpha_{n}=\gamma_{n}$ and $P_{n-1}$ is proportional to $Q_{n-1}$. Since both polynomials are monic, they are equal. ∎
Recall that we have relatively prime polynomials $N(u,s)$ and $D(u,s)$ such that
(62) $c(n,k+1)D(n,k)=c(n,k)N(n,k)\quad\forall n,k\in\mathbb{Z},$
and that we extend the domain of definition of the coefficients $c(n,k)$ to $\mathbb{Z}^{2}$ by setting $c(n,k)=0$ outside of the range $0\leq k\leq n$. We also have $c(n,n)=1$ for all $n\in\mathbb{Z}_{\geq 0}$.
Proof of Lemma 3.2.
Substituting $n=k$ in (62) we get $0=N(n,n)$. Thus the polynomial $N(u,u)$ vanishes identically, and thus $N$ is divisible by $(u-s)$. Substituting $k=-1$, we have $c(n,0)D(n,-1)=0$. We want to show that $D(n,-1)$ vanishes identically. Suppose the contrary. Then there exists a finite set $Y$ such that $c(n,0)=0$ for all $n\in\mathbb{Z}_{\geq 0}\smallsetminus Y$ and $c(n,0)\neq 0$ for all $n\in Y$. The set $Y$ is non-empty since $c(0,0)=0$. Let $n=\max(Y)+1$. Then from the three-term recursion we get $0=c(n+1,0)+\alpha_{n}c(n,0)+\beta_{n}c(n-1,0)=\beta_{n}c(n-1,0)$. Since $\beta_{n}\neq 0$, we get $c(n-1,0)=0$, contradicting $n-1\in Y$. Thus $D(u,-1)\equiv 0$, i.e. $D(u,s)$ is divisible by $s+1$. ∎
Proof of Lemma 3.3.
Since $c(n,n)=1$, (62) implies $c(n,n-1)N(n,n-1)=D(n,n-1)$. If $N(n,n-1)\equiv 0$ then $u-s-1$ divides $N$. Then $D(n,n-1)\equiv 0$ and thus $u-s-1$ divides also $D$. Since $N$ and $D$ are coprime, this is impossible, thus $N(n,n-1)\not\equiv 0$. Thus for $n$ outside a finite set we have
(63) $c(n,n-1)=\frac{D(n,n-1)}{N(n,n-1)},$
which is a well-defined rational function. Looking at the coefficient of $z^{k}$ in (16) we have
(64) $c(n,k-1)=c(n+1,k)+\alpha_{n}c(n,k)+\beta_{n}c(n-1,k)\quad\forall n,k\in\mathbb{Z}$
Substituting $k=n$ we obtain $\alpha_{n}=c(n,n-1)-c(n+1,n)$. By (63) this is an almost rational function on $\mathbb{Z}_{>0}$, i.e. coincides with a rational function outside a finite (possibly empty) set.
Substituting $k=n-1$ in (64) we have
(65) $c(n,n-2)=c(n+1,n-1)+\alpha_{n}c(n,n-1)+\beta_{n}\quad\forall n\in\mathbb{Z}$
By (62) we have
(66) ${N(n,n-2)}c(n,n-2)={D(n,n-2)}c(n,n-1)$
[…]
We just showed $m\in Y$. For any $l\in Y$ we have
[…]
$c(n,n-l-1)D(n,n-l-2)=d_{l}N(n,n-l-2).$
[…]
Now, if $m=1$ then taking $l:=m$ we obtain that $\beta_{n}$ is an almost rational function. Otherwise, $m\geq 3$, and $m-2\in Y$. Thus $c(n,n-m+1)$ is an almost rational function, and so is its shift $c(n-1,n-m)$. Since $m-1\notin Y$, this almost rational function is not almost zero. Thus
[…]
is an almost rational function.
Altogether, we showed that in all cases, both $\alpha_{n}$ and $\beta_{n}$ coincide with rational functions for all but finitely many $n$. ∎
Proof of Lemma 3.4.
[…]
(70) $c(n,n-m-2)=c(n+1,n-m-1)+\alpha_{n}c(n,n-m-1)+\beta_{n}c(n-1,n-m-1)$
[…]
Let us now show that $Sw=fw$. By (62) we have the equality $N(u,s)Sc=D(u,s)c$, which translates to the equality $N(u,s)Sw=D(u,s)w$ in $M$. This implies $Sw=fw$. ∎
A.1. Lemmas for families of Jacobi type
Proof of Lemma 6.1.
(i) We have to show that $D(s)$ has no zeros in $\mathbb{Z}_{\geq 0}$. Suppose the contrary, and let $k_{0}\in\mathbb{Z}_{\geq 0}$ be the largest integer zero. Let $n>k_{0}\in\mathbb{Z}_{\geq 0}$ be such that $N(n,i)\neq 0$ for any $0\leq i\leq k_{0}$. Let us show that for any such $n$ we have $c(n,k_{0})\neq 0$. Indeed, for any integer $i\in[k_{0},n]$ we have $c(n,i)\neq 0$. This follows by descending induction on $i$ from $c(n,n)=1$, $D(i+1)\neq 0$ and (62). Substituting $n$ and $k_{0}$ into (62) we obtain $N(n,k_{0})=0$. Since this holds for infinitely many $n$, we obtain that $N(u,k_{0})\equiv 0$, thus $N(u,s)$ is divisible by $s-k_{0}$, contradicting being coprime with $D(s)$.
(ii) follows similarly from (62) by descending induction on $k$, using $c(n,n)=1$ and $D(k)\neq 0$.
(iii) follows from (62) by descending induction on $k\in[0,n]$, the base case being $c(n,n)=1$. ∎
Proof of Lemma 6.2.
By (64) we have
$c(n,n-1)=c(n+1,n)+\alpha_{n}$
By Lemma 6.1 this implies
(71) $\alpha_{n}=\frac{c(n,n-1)-c(n+1,n)}{c(n,n)}=f(n,n-1)^{-1}-f(n+1,n)^{-1}$
for every $n>0$. Since this is a rational function, it equals $\alpha(n)$. Substituting $k={n-1}$ in (64) we have $c(n,n-2)=c(n+1,n-1)+\alpha_{n}c(n,n-1)+\beta_{n}$. This, together with the formula for $c(n,n-2)$ given by Lemma 6.1, implies that for every $n>1$ we have
(72) $\beta_{n}=c(n,n-2)-c(n+1,n-1)-\alpha_{n}c(n,n-1)=\\ f(n,n-2)^{-1}f(n,n-1)^{-1}-f(n+1,n-1)^{-1}f(n+1,n)^{-1}-f(n,n-1)^{-2}+f(n,n-1)^{-1}f(n+1,n)^{-1}$
(46). Since this is a rational function, it equals $\beta(n)$. Now, (45,46) follow from (72,71).
In (72), we used in particular that $c(n,n-2)=f(n,n-2)^{-1}f(n,n-1)^{-1}$ for all $n>1$. For $n=1$, we have $c(n,n-2)=c(1,-1)=0$. Thus $\beta_{1}=\beta(1)$ if and only if $f(1,-1)^{-1}f(1,0)^{-1}=0$. Since we already have the formulas for $f$ for all families of Jacobi type, we can check this case by case. The answer is that it holds for all families, except the Bessel family with $a=0$. ∎
Appendix B Examples of families of rational type
All the examples of rational families not of Jacobi type that we found are built in the same way. We find two families $P_{n},Q_{n}$ of Jacobi type that define the same module $M$, and have the same coefficients $\alpha_{n},\beta_{n}$ in the 3-term recursion, but $P_{1}\neq Q_{1}$. Then we define a new family $R_{n}^{\lambda}:=\mu P_{n}+(1-\mu)Q_{n}$. It lies in the same module and satisfies the same 3-term recursion, and thus is a rational family. In the two subsections below we do this for the Jacobi family with very special parameters, and for the two new families $E_{n}^{(c)}$ and $F_{n}^{(c)}$ (that actually define the same module). In §B.1.1 we consider a bit more elaborate construction: we let $R_{n}:=P_{n+1}-Q_{n+1}$ (which has degree $\leq n$ if $P_{n},Q_{n}$ are monic). In §B.3 below we summarize the examples in a table.
B.1. Jacobi polynomials with $a=1,b=l+1/2$
Consider the Jacobi family for $a=1$. Then
$\alpha=1/2,\text{ and }\beta=\frac{u^{2}-(b-1)^{2}}{4u^{2}-1}.$
We see that for $b^{\prime}=2-b$ we have the same recurrence relation. However, the 2 families are different, since they have different starting condition: $p_{1}=x-b/2$. Thus, any linear combination of the two families is quasi-orthogonal (unless $b$ is an integer). Their average, that has the initial condition $p_{1}=x-1/2$, is called “exceptional Jacobi polynomial’ - see [Ism09, Ch. 4, Remark 4.2.1]. It is not always a rational family, unless $b$ is a half-integer. If $b$ is half-integer then any linear combination is a rational family.
Let $P_{n}$ denote the Jacobi family for the parameters $a=1,b=l+1/2$ where $l\in\mathbb{Z}_{>0}$, and $Q_{n}$ denote the Jacobi family for the parameters $a=1,b^{\prime}=2-b=3/2-l$. Then $Q_{n,k}=\frac{r(k)}{r(n)}P_{n,k}$, where
(73) $r(t)=(t-l+3/2)_{(2l-1)}$
Consider the family $R^{l,\lambda}_{n}:=(\lambda P_{n}+Q_{n})/(\lambda+1)$ for $\lambda\neq-1$. Up to normalization, it is $g\mathcal{P}$, where
(74) $g(u,s)=\lambda+\frac{r(s)}{r(u)}=\frac{r(s)+\lambda r(u)}{r(u)}$
Consider first the case $l=1$. Then $r(t)=t+1/2$, and $g=\frac{\lambda(u+1/2)+s+1/2}{u+1/2}$. Thus
(75) $R^{1,\lambda}_{n}=\,_{3}F_{2}(-n,n+1,\lambda n+\frac{\lambda+3}{2};3/2,\lambda n+\frac{\lambda+1}{2};z)$
Thus this family satisfies a third order differential equation.
For $l>1$, the resulting $g$ does not seem to decompose to a product of linear factors. For $\lambda=1$ and $l=2$, the numerator of $g$ is $(s+u+1)(4s^{2}-4su+2s+4u^{2}+2u-3)$.
Actually, all the Jacobi families with $a=1,b=l+1/2$ lie in the same 1-dimensional module.
B.1.1. The case $\lambda=-1$
The polynomial $Q_{n}-R_{n}$ will have degree less than $n$. Thus let us define
(76) $R^{l}_{n}:=U(Q_{n}-P_{n})=\frac{r(s)-r(u+1)}{r(u+1)}P_{n+1}$
For this family we have
(77) $f=\frac{r(s+1)-r(u+1)}{r(s)-r(u+1)}\cdot\frac{(s-u-1)(s+u+2)}{(s+1)(s+l+1/2)}$
For $l=1$ we get that this is a Jacobi family with parameters $a=2,\,b=3/2$ in our notation. For $l=2$ we get a family of rational type that is not of Jacobi type.
B.2. The new families $E_{n}^{(c)}$, $F_{n}^{(c)}$
Let $c,\lambda\in\mathbb{C}$ with $c\notin\mathbb{Z}_{\leq 0}$, and define
(78) $g=(u+s+1)(s-u-c)-\lambda(u+s+c+1)(s-u),\\ p(t):=t(t-c),\,\,q(t):=(t+1)(t+c+1),\,\,w(s):=(s+1)(s+3/2)$
and
(79) $f(u,s):=\frac{p(s-u)q(u+s)g(u,s+1)}{w(s)g(u,s)},\,\,h=\frac{w(u)q(u+s)g(u,u)g(u+1,s)}{q(2u)q(2u+1)p(s-u-1)g(u+1,u+1)g(u,s)}$
Let $\mathcal{P}^{c,\lambda}$ be the polynomial family defined by $f$ via the formula $c(n,k)=\prod_{i=k}^{n-1}f(n,i)^{-1}$. Let
(80) $\alpha=-\frac{1}{2(2u+c)(2u+c+2)}\text{ and }\beta=\frac{1}{16(2u+c+1)(2u+c)^{2}(2u+c-1)}$
Then by a straightforward computation we get $c_{n,k+1}=f(n,k)c_{n,k},\,\,c_{n+1,k}=h(n,k)c_{n,k},$ and
(81) $S^{-1}(f)^{-1}=h+\alpha+\beta/U^{-1}(h).$
Thus, $zP^{c,\lambda}_{n}=P^{c,\lambda}_{n+1}+\alpha(n)P^{c,\lambda}_{n}+\beta(n)P^{c,\lambda}_{n-1}$ and thus the family is quasi-orthogonal. Note that for $\lambda=0$ we obtain the family $F^{(c)}_{n}$ and for $\lambda=1$ we obtain the family $E_{n}^{(c+1)}$ this can be seen by comparing (79) with (40). The $\alpha$ and $\beta$ do not depend on $\lambda$ and are the same as for $F_{n}^{(c)}$ and $E_{n}^{(c+1)}$. For $\lambda\notin\{0,1\}$ the family is of rational type but not of Jacobi type, since the denominator of $f$ depends on $u$.
B.3. Summarizing table
The three families described in this subsection can be presented up to renormalization as $g(n,z\partial_{z})Q_{n},$ where $g$ and $Q_{n}$ are given by the following table.
$Q_{n}$ $g(u,s)$
Jacobi with $a=1$, $b=l+1/2,\,l\in\mathbb{Z}_{>0}$ $(s-l+3/2)_{(2l-1)}+\lambda(u-l+3/2)_{(2l-1)},\,\lambda\neq-1$
Jacobi with $a=2$, $b=l+1/2,\,l\in\mathbb{Z}_{>0}$ $\frac{(s-l+3/2)_{(2l-1)}-(u-l+5/2)_{(2l-1)}}{s-u-1},\,\lambda\neq-1$
$\,{}_{4}F_{1}(-n,-n-c,n+1,n+c+1;3/2;z)$ $(u+s+1)(s-u-c)-\lambda(u+s+c+1)(s-u)$
Acknowledgements
It is a real pleasure to thank Ehud de Shalit and Don Zagier for fruitful discussions, and Tom Koornwinder and Mourad Ismail for their insightful and helpful remarks on a preliminary version of the paper. The research of J. Bernstein was partially supported by the ISF grant 1400/19. The research of S. Sahi was partially supported by NSF grants DMS-1939600 and 2001537, and Simons Foundation grant 509766. The research of D. Gourevitch was partially supported by BSF grant 2019724 and ISF grant 1781/23. A significant part of the research was done during the visits of S. Sahi at the Weizmann Institute, the visit of D. Gourevitch at the IAS, and the visit of J. Bernstein to the MPIM, Bonn in 2023. We thank the three institutes for their hospitality, and the BSF for funding two of these visits.